OpenStudy (anonymous):

Divide (-3+5i)/(5-4i)

OpenStudy (anonymous):

\[\frac{-3+5i}{5-4i}\times \frac{5+4i}{5+4i}=\frac{(-3+5i)(5+4i)}{5^2+4^2}\] is a start

OpenStudy (anonymous):

working..